OpenStudy (anonymous):

Best answer gets a medal! solve the following using the quadratic formula (answer in simplest exact values) 5x^2-11=0

OpenStudy (anonymous):

\[\sqrt{\frac{ 55 }{ ? }}\]

OpenStudy (anonymous):

\[\sqrt{\frac{ 55 }{3 }}\]

OpenStudy (anonymous):

\[\sqrt{220}\div10\]